Find the area of triangle determinied by the point P(-3,1-1) Q(-2,0,5) and R(-4,-1,2).
1
Expert's answer
2020-10-26T19:00:00-0400

Firstly we should find vectors "\overline{PQ}" and "\overline{QR}" :

"\overline{PQ}=(1,-1,6)"

"\overline{QR}" "=(-2,-1,-3)"

The area of ​​a triangle PQR is half the area of ​​a parallelogram built on vectors "\overline{PQ}" and "\overline{QR}". The area of ​​a parallelogram built on vectors "\overline{PQ}" and "\overline{QR}" is the modulus of the vector product "\overline{PQ}" and "\overline{QR}", and therefore the area of ​​triangle PQR is:

"S_{PQR}=\frac{1}{2}|\overline{PQ}\times\overline{QR}|"

"\overline{PQ}\times\overline{QR}=" 9*"\overline{i}" -9*"\overline{j}" -3*"\overline{k}"

"S_{PQR}=\frac{1}{2}\sqrt{81+81+9}=\frac{3\sqrt{19}}{2}"